# Order isomorphisms of sup-stable function spaces: continuous, Lipschitz, c-convex, and beyond

### v1

@@ 2.2 The $e_{x}$, the Dirac-like inf-irreducible functions of $\mathcal{G}$
$[f]:=\{g\in\mathcal{G}\,|\,\exists\alpha\,\in\mathbb{R},\,f-\alpha\leq g\leq f+\alpha\}$ (15)
@@
Let $x\in\mathscr{X}$. Since $\mathcal{G}$ is proper, we can fix $u_{x}\in\mathcal{G}$ such that $u_{x}(x)\in\mathbb{R}$. So $e_{x}\geq u_{x}-u_{x}(x)$, whence $e_{x}(x)=0$. We have that $e_{x}\in\mathcal{G}$ as $\mathcal{G}$ is sup-stable. If $e_{x}=\inf(f,g)$ for some $f,g\in\mathcal{G}$ then w.l.o.g. we can assume that $f(x)=e_{x}(x)=0$ so, by maximality of $e_{x}$, $f\leq e_{x}=\inf(f,g)$ and $f=e_{x}$. Take $f\in\mathcal{G}$ with $[f]$ maximal. Since $f\leq e_{x}+f(x)$ for all $x\in\operatorname*{Dom}(f)$ i.e. $[e_{x}]\geq[f]$, we have that $[e_{x}]\leq[f]$ implies that $[e_{x}]=[f]$. Fix $x\neq x^{\prime}$ and $\lambda\in\overline{\mathbb{R}}$ and $g_{1},g_{2}\in\mathcal{G}$ as in 1. Assume that $e_{x}=e_{x^{\prime}}+\lambda$. We can already exclude $\lambda\in\{\pm\infty\}$ since $e_{x}$ has a finite value in $x$. Moreover by definition of $e_{x}$ we have
@@
Evaluating in $x^{\prime}$ and $x$ the above expressions gives $\lambda\geq g_{1}(x^{\prime})-g_{1}(x)\geq\lambda$. The same expression holds for $g_{2}$, which contradicts 1. So $e_{x}\neq e_{x^{\prime}}+\lambda$. Eq.(17) is just a rewriting of (16).
@@
Consider the set $\mathcal{G}$ of increasing functions from $\mathbb{R}$ to $\overline{\mathbb{R}}$. It is sup and inf-stable. A quick computation gives $e_{x}=\delta^{\top}_{\{y\in\mathbb{R}|y\leq x\}}$ where $\delta^{\top}_{A}$ denotes the indicator function of a subset $A\subset\mathbb{R}$. This $e_{x}$ looks like a Dirac function, but $[e_{x}]$ is strictly smaller than $[e_{x^{\prime}}]$ for any $x^{\prime}<x$, so no Archimedean class is maximal.
@@
$\overline{\mathcal{G}}^{\inf}=\{\inf_{x}e_{x}(\cdot)+w_{x}\,|\,w_{x}\in\overline{\mathbb{R}}\}\text{ and }e_{x}(\cdot)=\sup\{u\in\overline{\mathcal{G}}^{\inf}\,|\,u(x)<0\}$ (19)
@@
The first statement is just a consequence of the fact that $\mathcal{G}\subset\overline{\mathcal{G}}^{\inf}$. Since $e_{x}(\cdot)+w_{x}\in\mathcal{G}$, by definition $\{\inf_{x}e_{x}(\cdot)+w_{x}\,|\,w_{x}\in\overline{\mathbb{R}}\}\subset\overline{\mathcal{G}}^{\inf}$. Take $f\in\overline{\mathcal{G}}^{\inf}$, so $f=\inf_{\alpha\in\mathcal{A}}h_{\alpha}$ for $\mathcal{A}$ an index set and $(h_{\alpha})_{\alpha\in\mathcal{A}}\subset\mathcal{G}$. By (18), each $h_{\alpha}$ has a representation by the $e_{x}$, commute the infima and set $w_{x}=\inf_{\alpha\in\mathcal{A}}h_{\alpha}(x)$ to conclude that $\overline{\mathcal{G}}^{\inf}=\{\inf_{x}e_{x}(\cdot)+w_{x}\,|\,w_{x}\in\overline{\mathbb{R}}\}$.
Set $\tilde{e}_{x}^{<}(\cdot)=\sup\{u\in\overline{\mathcal{G}}^{\inf}\,|\,u(x)<0\}$ and $\tilde{e}_{x}(\cdot)=\sup\{u\in\overline{\mathcal{G}}^{\inf}\,|\,u(x)\leq 0\}$. By definition, $\tilde{e}_{x}^{<}\leq\tilde{e}_{x}$, and, for all $\epsilon>0$, $\tilde{e}_{x}-\epsilon\leq\tilde{e}_{x}^{<}$. Taking the limit $\epsilon\to 0^{+}$, $\tilde{e}_{x}^{<}=\tilde{e}_{x}$. Since $\mathcal{G}\subset\overline{\mathcal{G}}^{\inf}$, $e_{x}\leq\tilde{e}_{x}$. Let $u=\inf_{\alpha\in\mathcal{A}}h_{\alpha}\in\overline{\mathcal{G}}^{\inf}$ such that $u(x)<0$, fix $\alpha\in\mathcal{A}$ such that $h_{\alpha}(x)\leq 0$. As $h_{\alpha}\in\mathcal{G}$, we have $u\leq h_{\alpha}\leq e_{x}$. Take the supremum over such $u$ to conclude that $\tilde{e}_{x}^{<}\leq e_{x}$, which concludes the proof. ∎
For sets $\mathcal{G}=\operatorname*{Rg}(B)$ we simply have that the $e_{x}$ are the image of the $b(x,\cdot)$.
@@
Take first $u=b(\cdot,y)$ to show that $\bar{B}e_{x}(z)$ is smaller than $b(x,y)$ and then $z=x$ to show it is also larger. Hence $\bar{B}e_{x}(y)=b(x,y)$. And conversely since $\bar{B}\bar{B}^{\circ}\bar{B}=\bar{B}$, $\bar{B}^{\circ}$ is a (max,+) anti-isomorphism over $\mathcal{G}$ and $e_{x}=\bar{B}b(x,\cdot)$. ∎
Using (17), taking $\mathcal{G}$ to be the set of 1-Lipschitz functions over a metric space $(\mathscr{X},d)$ leads to $e_{x}=d(\cdot,x)$, and, more generally, for $(C,p)$-Hölder functions, $e_{x}=Cd(\cdot,x)^{p}$ for $C>0$ and $p>0$. When $\mathcal{G}$ is the set of (convex) lower semicontinuous functions over a topological space $\mathscr{X}$, we have $e_{x}=\delta_{x}^{\top}$. Actually, whenever $(\delta_{x}^{\top})_{x\in\mathscr{X}}\subset\mathcal{G}$, the Dirac masses are necessarily the $e_{x}$. There are many spaces for which it is the case, for instance all those that are defined through a growth condition, see [Sin97, p126]. We now turn to our main result.
@@
    there exists $g:\mathscr{X}\rightarrow\mathbb{R}$ and $\phi:\mathscr{X}\rightarrow\mathscr{Y}$ such that
@@
    and $\phi$ is bijective and, for all $f\in\mathscr{F}$, $h\in\mathcal{G}$, $(g+f\circ\phi)\in\mathcal{G}$ and $(-g\circ\phi^{-1}+h\circ\phi^{-1})\in\mathscr{F}$;
@@
Conversely every $J$ satisfying ii) is order preserving and maps $\mathscr{F}$ to $\mathcal{G}$, and has an inverse with the same properties, so it is a (max,+)-isomorphism.
@@
Fix $y\in\mathscr{Y}$ and recall that $e^{\prime}_{y}(\cdot):=\sup\{v\in\mathscr{F}\,|\,v(y)\leq 0\}$. By assumption i), $Je^{\prime}_{y}=e_{\psi(y)}+\bar{g}(y)$ with $\bar{g}(y)=Je^{\prime}_{y}(\psi(y))\in\mathbb{R}$. Conversely there exists a $\varphi:\mathscr{X}\to\mathscr{Y}$ and $\lambda_{x}\in\mathbb{R}$ such that $e_{x}=J(e_{\varphi(x)}+\lambda_{x})=e_{\psi(\varphi(x))}+\lambda_{x}+\bar{g}(\varphi(x))$. However from 17, we know that 1 then ensures that $x=\psi(\varphi(x))$, so $\psi$ is bijective.
@@
$=\lambda^{\prime}+g(\phi^{-1}(y))+\sup\{u\in\mathcal{G}\,|\,u(\phi^{-1}(y))\leq 0\}=\lambda^{\prime}+g(\phi^{-1}(y))+e_{\phi^{-1}(y)}.$
@@
Figure 1: The isomorphic spaces $\operatorname*{Rg}_{\mathbb{R}}(A)$ and $\operatorname*{Rg}_{\mathbb{R}}(B)$ (here intersected with $x_{1}+x_{2}+x_{3}=0$) do not have (max,+)-isomorphisms of the form (21).
@@
This example is illustrated in Figure 1. One can vizualize here that $\operatorname*{Rg}_{\mathbb{R}}(A)$ and $\operatorname*{Rg}_{\mathbb{R}}(B)$ have the same diameters in Hilbert’s seminorm (the diameter is achieved by the distance between the two first columns, in each case). It is not difficult to check that the diameter is the only invariant for (max,+) spaces with two finite generators belonging to some $\mathbb{R}^{n}$ meaning that two such spaces are isomorphic iff they have the same diameter, which explains how this example was constructed. Note also that $A_{\cdot 1},A_{\cdot 2}$ are sup-irreducible and relatively-inf-irreducible, but that $A_{\cdot 3}$ is neither, even though $\mathbb{R}e+\{A_{\cdot 1},A_{\cdot 2},A_{\cdot 3}\}$ are the only inf-irreducible elements in $\operatorname*{Rg}(A)$. This underlines that the notions of irreducibility are distinct.
However when the Dirac masses belong to the sets, every (max,+)-isomorphism is affine.
@@
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ (resp. ¯$\overline{\mathbb{R}}^{\mathscr{Y}}$) for some sets $\mathscr{X}$ and $\mathscr{Y}$, with both $\mathscr{F}$ and $\mathcal{G}$ satisfying 1. Let $J$ be a (max,+)-isomorphism from $\mathscr{F}$ onto $\mathcal{G}$. Assume that $(\delta^{\top}_{y})_{y\in\mathscr{Y}}\subset\mathscr{F}$ and $(\delta^{\top}_{x})_{x\in\mathscr{X}}\subset\mathcal{G}$. Then the statements of 21 hold, in particular there exists $g:\mathscr{X}\to\mathbb{R}$ and a bijective $\phi:\mathscr{X}\to\mathscr{Y}$ such that $Jf=g+f\circ\phi$.
@@
Fix $x\in\mathscr{X}$ and $y\in\mathscr{Y}$, $\lambda\in\mathbb{R}$ and $\lambda^{\prime}\in\mathbb{R}$. By definition (16), we have that $e_{x}=\delta^{\top}_{x}$ and $e^{\prime}_{y}=\delta^{\top}_{y}$. Moreover $[\delta^{\top}_{x}+\lambda]$ is maximal since if there exists $h\in\mathcal{G}$ and $\alpha\in\mathbb{R}$ such that $\delta^{\top}_{x}+\lambda\leq h+\alpha$, then $h=\delta^{\top}_{x}+h(x)$, similarly $[\delta^{\top}_{y}+\lambda^{\prime}]$ is maximal. Moreover 17 guarantees that these are the only one maximal Archimedean classes, and 16 shows that $J$ sends an $f\in\mathscr{F}$ with $[f]$ maximal onto a $Jf\in\mathcal{G}$ with $[Jf]$ maximal. Hence we “send Dirac masses onto Dirac masses” and 21 i) holds.
@@
We have shown in [AFG23, Proposition 6.2] that complete subspaces $\mathcal{G}\subset\mathbb{R}^{\mathscr{X}}$ that are both inf and sup-stable can always be described as the set of 1-Lipschitz functions for a weak $\overline{\mathbb{R}}$-valued metric $\tilde{d}:\mathscr{X}\times\mathscr{X}\to\overline{\mathbb{R}}$ and $\tilde{d}(z,x)=e_{x}(z)$. For inf-stable sets $\mathcal{G}$, the $e_{x}$ are (relatively)-inf-irreducible functions. It is possible sometimes to show they are the only ones even without resorting to Archimedean classes. For instance, for the set $\mathcal{G}_{\nearrow}^{\mathbb{R}}$ of increasing functions from $\overline{\mathbb{R}}$ to $\overline{\mathbb{R}}$, consider a candidate inf-irreducible $f$ and take $x_{0}\in\operatorname*{Dom}(f)$, $A_{0}=\{x\in\mathbb{R}\,|\,f(x)>f(x_{0})\}$, then $f=\inf[\sup(f(x_{0}),f),f+i_{A_{0}}]$. So 21-i) holds for all (max,+)-isomorphisms $J:\mathcal{G}_{\nearrow}^{\mathbb{R}}\to\mathcal{G}_{\nearrow}^{\mathbb{R}}$ since they preserve inf-irreducibility. This is a case where there are no maximal Archimedean classes as discussed in 18. As we will see in Section 3.2 and hinted at already in 17, maximal Archimedean classes when they exist allow to separate within inf-irreducible points the $e_{x}$ from the horoboundary elements, among which the Busemann points.
2.3 Sup-irreducible functions and a Krein-Milman theorem
We have explored the inf-irreducible elements of a sup-stable $\mathcal{G}$ by studying the $e_{x}$, can we symmetrize the arguments to obtain the sup-irreducible elements? In general not, because $\mathcal{G}$ has no reason to be inf-stable. Nevertheless we can use the notion of minimal element of a subset $\mathcal{A}\subset\mathcal{G}$, i.e. the $u\in\mathcal{A}$ such that for any $v\in\mathcal{A}$ with $v\leq u$, we have $v=u$. This gives us a Krein–Milman theorem

### v2

@@ 2.2 The $e_{x}$, the Dirac-like inf-irreducible functions of $\mathcal{G}$
$[f]:=\{g\in\mathcal{G}\,|\,\exists\alpha\,\in\mathbb{R},\,f-\alpha\leq g\leq f+\alpha\}.$ (17)
@@
Let $x\in\mathscr{X}$. Since $\mathcal{G}$ is proper, we can fix $u_{x}\in\mathcal{G}$ such that $u_{x}(x)\in\mathbb{R}$. So $e_{x}\geq u_{x}-u_{x}(x)$, whence $e_{x}(x)\geq 0$, while by definition $e_{x}(x)\leq 0$. We have that $e_{x}\in\mathcal{G}$ as $\mathcal{G}$ is sup-stable. If $e_{x}=\inf(f,g)$ for some $f,g\in\mathcal{G}$ then w.l.o.g. we can assume that $f(x)=e_{x}(x)=0$ so, by maximality of $e_{x}$, $f\leq e_{x}=\inf(f,g)$ and $f=e_{x}$. Take $f\in\mathcal{G}$ with $[f]$ maximal. Since $f\leq e_{x}+f(x)$ for all $x\in\operatorname*{Dom}(f)$ i.e. $[e_{x}]\geq[f]$, we have that $[e_{x}]\leq[f]$ implies that $[e_{x}]=[f]$. Fix $x\neq x^{\prime}$ and $\lambda\in\overline{\mathbb{R}}$ and $g_{1},g_{2}\in\mathcal{G}$ as in 1. Assume that $e_{x}=e_{x^{\prime}}+\lambda$. We can already exclude $\lambda\in\{\pm\infty\}$ since $e_{x}$ has a finite value in $x$. Moreover by definition of $e_{x}$ we have
@@
Evaluating in $x^{\prime}$ and $x$ the above expressions gives $\lambda\geq g_{1}(x^{\prime})-g_{1}(x)\geq\lambda$. The same holds for $g_{2}$, which contradicts 1. So $e_{x}\neq e_{x^{\prime}}+\lambda$. Eq.(19) is just a rewriting of (18).
@@
Consider the set $\mathcal{G}$ of increasing functions from $\mathbb{R}$ to $\overline{\mathbb{R}}$. It is sup and inf-stable. A quick computation gives $e_{x}=\delta^{\top}_{\{y\in\mathbb{R}|y\leq x\}}$ where $\delta^{\top}_{A}$ denotes the indicator function of a subset $A\subset\mathbb{R}$. This $e_{x}$ looks like a Dirac function, but $[e_{x}]$ is strictly smaller than $[e_{x^{\prime}}]$ for any $x^{\prime}<x$, so no Archimedean class is maximal. Moreover the constant functions are inf-irreducible, but do not belong to $\{e_{x}+\lambda\}_{\lambda\in\mathbb{R},x\in\mathbb{R}}$.
@@
$\overline{\mathcal{G}}^{\inf}=\{\inf_{x}e_{x}(\cdot)+w_{x}\,|\,w_{x}\in\overline{\mathbb{R}}\}\text{ and }e_{x}(\cdot)=\sup\{u\in\overline{\mathcal{G}}^{\inf}\,|\,u(x)<0\}.$ (21)
@@
The first statement is just a consequence of the fact that $\mathcal{G}\subset\overline{\mathcal{G}}^{\inf}$. Concerning (21), since $e_{x}(\cdot)+w_{x}\in\mathcal{G}$, by definition $\{\inf_{x}e_{x}(\cdot)+w_{x}\,|\,w_{x}\in\overline{\mathbb{R}}\}\subset\overline{\mathcal{G}}^{\inf}$. Take $f\in\overline{\mathcal{G}}^{\inf}$, so $f=\inf_{\alpha\in\mathcal{A}}h_{\alpha}$ for $\mathcal{A}$ an index set and $(h_{\alpha})_{\alpha\in\mathcal{A}}\subset\mathcal{G}$. By (20), each $h_{\alpha}$ has a representation by the $e_{x}$, commute the infima and set $w_{x}=\inf_{\alpha\in\mathcal{A}}h_{\alpha}(x)$ to conclude that $\overline{\mathcal{G}}^{\inf}=\{\inf_{x}e_{x}(\cdot)+w_{x}\,|\,w_{x}\in\overline{\mathbb{R}}\}$.
Set $\tilde{e}_{x}^{<}(\cdot)=\sup\{u\in\overline{\mathcal{G}}^{\inf}\,|\,u(x)<0\}$ and $\tilde{e}_{x}(\cdot)=\sup\{u\in\overline{\mathcal{G}}^{\inf}\,|\,u(x)\leq 0\}$. By definition, $\tilde{e}_{x}^{<}\leq\tilde{e}_{x}$, and, for all $\epsilon>0$, $\tilde{e}_{x}-\epsilon\leq\tilde{e}_{x}^{<}$. Taking the limit $\epsilon\to 0^{+}$, $\tilde{e}_{x}^{<}=\tilde{e}_{x}$. Since $\mathcal{G}\subset\overline{\mathcal{G}}^{\inf}$, $e_{x}\leq\tilde{e}_{x}$. Let $u=\inf_{\alpha\in\mathcal{A}}h_{\alpha}\in\overline{\mathcal{G}}^{\inf}$ such that $u(x)<0$ and fix $\alpha\in\mathcal{A}$ such that $h_{\alpha}(x)\leq 0$. As $h_{\alpha}\in\mathcal{G}$, we have $u\leq h_{\alpha}\leq e_{x}$. Taking the supremum over such $u$ entails that $\tilde{e}_{x}^{<}\leq e_{x}$, which concludes the proof. ∎
For sets $\mathcal{G}=\operatorname*{Rg}(B)$ we simply have that $e_{x}$ is the image of $b(x,\cdot)$.
@@
Take first $u=b(\cdot,y)$ to show that $\bar{B}^{\circ}e_{x}(z)$ is smaller than $b(x,y)$ and then $z=x$ to show it is also larger. Hence $\bar{B}^{\circ}e_{x}(y)=b(x,y)$. And conversely since $\bar{B}\bar{B}^{\circ}\bar{B}=\bar{B}$, $\bar{B}^{\circ}$ is a (max,+) anti-isomorphism over $\mathcal{G}$ and $e_{x}=\bar{B}b(x,\cdot)$. ∎
Using (19), taking $\mathcal{G}$ to be the set of 1-Lipschitz functions over a metric space $(\mathscr{X},d)$ leads to $e_{x}=d(\cdot,x)$, and, more generally, for $(C,p)$-Hölder functions, $e_{x}=Cd(\cdot,x)^{p}$ for $C>0$ and $p>0$. When $\mathcal{G}$ is the set of (convex) lower semicontinuous functions over a topological space $\mathscr{X}$, we have $e_{x}=\delta_{x}^{\top}$. Actually, whenever $\{\delta_{x}^{\top}\}_{x\in\mathscr{X}}\subset\mathcal{G}$, the Dirac masses are necessarily the $e_{x}$. There are many spaces for which it is the case, for instance all those that are defined through a growth condition, see [Sin97, p127]. We now turn to our first main result.
@@
    there exists a function $g:\mathscr{X}\rightarrow\mathbb{R}$ and a bijection $\phi:\mathscr{X}\rightarrow\mathscr{Y}$ such that
@@
    and, for all $f\in\mathscr{F}$, $h\in\mathcal{G}$, $(g+f\circ\phi)\in\mathcal{G}$ and $(-g\circ\phi^{-1}+h\circ\phi^{-1})\in\mathscr{F}$;
@@
Conversely every $J$ satisfying ii) is order preserving, commutes with the addition of scalars, maps $\mathscr{F}$ to $\mathcal{G}$, and has an inverse with the same properties, so it is a (max,+)-isomorphism. Moreover the representation $(g,\phi)$ in (23) is unique.
The representation (23) is analogous to the polar decomposition of matrices in the general real linear group as a product of a symmetric positive matrix and an orthogonal matrix.
@@
Fix $y\in\mathscr{Y}$ and recall that $e^{\prime}_{y}(\cdot):=\sup\{v\in\mathscr{F}\,|\,v(y)\leq 0\}$. By assumption i), $Je^{\prime}_{y}=e_{\psi(y)}+\bar{g}(y)$ with $\bar{g}(y)=Je^{\prime}_{y}(\psi(y))\in\mathbb{R}$. Conversely there exists a $\varphi:\mathscr{X}\to\mathscr{Y}$ and $\lambda_{x}\in\mathbb{R}$ such that $e_{x}=J(e_{\varphi(x)}+\lambda_{x})=e_{\psi(\varphi(x))}+\lambda_{x}+\bar{g}(\varphi(x))$. However from 17, we know that 1 then ensures that $x=\psi(\varphi(x))$, so $\psi$ is surjective. Similarly, consider $Je^{\prime}_{y}=Je^{\prime}_{y^{\prime}}$, to get that $\psi$ is injective.
@@
$=\lambda^{\prime}+g(\phi^{-1}(y))+\sup\{u\in\mathcal{G}\,|\,u(\phi^{-1}(y))\leq 0\}=\lambda^{\prime}+g(\phi^{-1}(y))+e_{\phi^{-1}(y)},$
so i) holds. Assume that $Jf=g+f\circ\phi=\tilde{g}+f\circ\tilde{\phi}$ for some alternative $\tilde{g}$ and $\tilde{\phi}$ satisfying (23). Then the above computation gives that $g(\phi^{-1}(y))+e_{\phi^{-1}(y)}=\tilde{g}(\tilde{\phi}^{-1}(y))+e_{\tilde{\phi}^{-1}(y)}$ for all $y\in\mathscr{Y}$. 17 then gives that necessarily $\phi=\tilde{\phi}$, and, evaluating in $\phi^{-1}(y)$, $g=\tilde{g}$. So the representation (23) is unique.
@@
[…]
Figure 1: The isomorphic spaces $\operatorname*{Rg}_{\mathbb{R}}(A)$ and $\operatorname*{Rg}_{\mathbb{R}}(B)$ (here intersected with the hyperplane $x_{1}+x_{2}+x_{3}=0$) do not have (max,+)-isomorphisms of the form (23).
@@
This example is illustrated in Figure 1. One can vizualize here that $\operatorname*{Rg}_{\mathbb{R}}(A)$ and $\operatorname*{Rg}_{\mathbb{R}}(B)$ have the same diameters in Hilbert’s seminorm (the diameter is achieved by the distance between the two first columns, in each case). It is not difficult to check that the diameter is the only invariant for (max,+) spaces with two finite generators belonging to some $\mathbb{R}^{n}$, meaning that two such spaces are isomorphic iff they have the same diameter, which explains how this example was constructed. Note also that $A_{\cdot 1},A_{\cdot 2}$ are sup-irreducible and relatively-inf-irreducible, but that $A_{\cdot 3}=\inf^{\operatorname*{Rg}(A)}((1,0,1),(0,1,1))=\sup(A_{\cdot 1},A_{\cdot 2})$ is neither, even though $\mathbb{R}e+\{A_{\cdot 1},A_{\cdot 2},A_{\cdot 3}\}$ are the only inf-irreducible elements in $\operatorname*{Rg}(A)$. This underlines that the notions of irreducibility are distinct.
Nevertheless when the Dirac masses belong to the sets, every (max,+)-isomorphism is affine.
@@
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ (resp. ¯$\overline{\mathbb{R}}^{\mathscr{Y}}$) for some sets $\mathscr{X}$ and $\mathscr{Y}$, with both $\mathscr{F}$ and $\mathcal{G}$ satisfying 1. Let $J$ be a (max,+)-isomorphism from $\mathscr{F}$ onto $\mathcal{G}$. Assume that $\{\delta^{\top}_{y}\}_{y\in\mathscr{Y}}\subset\mathscr{F}$ and $\{\delta^{\top}_{x}\}_{x\in\mathscr{X}}\subset\mathcal{G}$. Then the statements of 21 hold, in particular there exists $g:\mathscr{X}\to\mathbb{R}$ and a bijective $\phi:\mathscr{X}\to\mathscr{Y}$ such that $Jf=g+f\circ\phi$.
@@
Fix $x\in\mathscr{X}$ and $y\in\mathscr{Y}$, $\lambda\in\mathbb{R}$ and $\lambda^{\prime}\in\mathbb{R}$. By definition (18), we have that $e_{x}=\delta^{\top}_{x}$ and $e^{\prime}_{y}=\delta^{\top}_{y}$. Moreover $[\delta^{\top}_{x}+\lambda]$ is maximal since if there exists $h\in\mathcal{G}$ and $\alpha\in\mathbb{R}$ such that $\delta^{\top}_{x}+\lambda\leq h+\alpha$, then $h=\delta^{\top}_{x}+h(x)$, similarly $[\delta^{\top}_{y}+\lambda^{\prime}]$ is maximal. Moreover 17 guarantees that these are the only maximal Archimedean classes, and 16 shows that $J$ sends an $f\in\mathscr{F}$ with $[f]$ maximal onto a $Jf\in\mathcal{G}$ with $[Jf]$ maximal. Hence we “send Dirac masses onto Dirac masses” and 21 i) holds.
@@
We have shown in [AFG24, Proposition 6.2] that complete subspaces $\mathcal{G}\subset\mathbb{R}^{\mathscr{X}}$ that are both inf and sup-stable can always be described as the set of 1-Lipschitz functions for a weak $\overline{\mathbb{R}}$-valued metric $\tilde{d}:\mathscr{X}\times\mathscr{X}\to\overline{\mathbb{R}}$ and $\tilde{d}(z,x)=e_{x}(z)$. For inf-stable sets $\mathcal{G}$, the $e_{x}$ are (relatively)-inf-irreducible functions. It is possible sometimes to show they are the only ones even without resorting to Archimedean classes. For instance, for the set $\mathcal{G}_{\nearrow}^{\mathbb{R}}$ of increasing functions from $\overline{\mathbb{R}}$ to $\overline{\mathbb{R}}$, consider a candidate inf-irreducible $f$ and take $x_{0}\in\operatorname*{Dom}(f)\cap\mathbb{R}$, $A_{0}=\{x\in\mathbb{R}\,|\,f(x)>f(x_{0})\}$, then $f=\inf[\sup(f(x_{0}),f),f+i_{A_{0}}]$, which forces $f=e_{x^{\prime}_{0}}+\lambda_{0}$ for some $x^{\prime}_{0}\in\overline{\mathbb{R}}$, and for $x^{\prime}_{0}=+\infty$ we recover the inf-irreducible constant functions, which are Busemann points. So 21-i) holds for all (max,+)-isomorphisms $J:\mathcal{G}_{\nearrow}^{\mathbb{R}}\to\mathcal{G}_{\nearrow}^{\mathbb{R}}$ since they preserve inf-irreducibility. This is a case where there are no maximal Archimedean classes as discussed in 18. As we will see in Section 3.2 and hinted at already in 17, maximal Archimedean classes when they exist allow to separate within inf-irreducible points the $e_{x}$ from the horoboundary elements, among which the Busemann points.
2.3 Sup-irreducible functions and a Krein–Milman theorem
We have explored the inf-irreducible elements of a sup-stable $\mathcal{G}$ by studying the $e_{x}$, can we symmetrize the arguments to obtain the sup-irreducible elements? In general not, because $\mathcal{G}$ has no reason to be inf-stable. Nevertheless we can use the notion of minimal element of a subset $\mathcal{A}\subset\mathcal{G}$, i.e. the $u\in\mathcal{A}$ such that for any $v\in\mathcal{A}$ with $v\leq u$, we have $v=u$. This gives us a Krein–Milman theorem.